ISS On-Orbit Status 02/11/10
February 11, 2010
All ISS systems continue to function nominally, except those noted previously or below. Sleep shift in effect: The ISS crew’s workday began last evening at 4:14pm and ended this morning at 7:44 EST (see time table at bottom).
At wake-up last evening, FE-1 Suraev began his day with the regular daily checkup of the aerosol filters at the Elektron O2
generator. [The filters were installed by Maxim on 10/19/09 in gaps between the BZh Liquid Unit and the oxygen outlet pipe (filter FA-K) plus hydrogen outlet pipe (filter FA-V). Photographs are to be taken if the filter packing is discolored.]
Also at wake-up, the FE-1 terminated his 9th
experiment session, started last night, for the long-term Russian sleep study MBI-12/SONOKARD, by taking the recording device from his SONOKARD sports shirt pocket and later copying the measurements to the RSE-MED laptop for subsequent downlink to the ground. [SONOKARD objectives are stated to (1) study the feasibility of obtaining the maximum of data through computer processing of records obtained overnight, (2) systematically record the crewmember’s physiological functions during sleep, (3) study the feasibility of obtaining real-time crew health data. Investigators believe that contactless acquisition of cardiorespiratory data over the night period could serve as a basis for developing efficient criteria for evaluating and predicting adaptive capability of human body in long-duration space flight.]
CDR Williams & FE-6 Creamer continued their new week-long session of the experiment SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight), TJ’s second, donning their Actiwatches, from which to log data to the HRF-1 (Human Research Facility 1) laptop. [To monitor the crewmembers’ sleep/wake patterns and light exposure, the crewmembers sometimes wear a special Actiwatch device which measures the light levels encountered by them as well as their patterns of sleep and activity throughout the Expedition and use the payload software for data logging and filling in questionnaire entries in the experiment’s laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days.]
Working several hours in the Lab on the WRS (Water Recovery System), CDR Williams first completed the installation of the new 20A-delivered UPA DA (Urine Processor Assembly / Distillation Assembly), then installed the new FCPA (Fluids Control Pump Assembly)
Afterwards, Jeff offloaded the WPA WWT (Water Processor Assembly / Waste Water Tank) and installed the (WPA) Pump Separator Filter. Later, the CDR filled the WSTA (Waste Storage Tank Assembly) to start UPA processing.
Suraev recharged the DZZ-13 battery (AIP-01), and then conducted a “Dark Current Recording” observation session with the Russian DZZ-13 RUSALKA (“Mermaid”) science experiment, to evaluate RUSALKA science hardware performance supported by ground specialist tagup. The equipment was then closed down and the data downlinked. [RUSALKA ops involve calibration and tests of research equipment relating to the Sun and the Earth's limb at sunset (atmosphere lighted). To be tested are the procedure for remote determination of Methane (CH4) & Carbon Dioxide (CO2) content in the atmosphere (in the First Phase), measurement of CH4 & CO2 content in the atmosphere and reception of data on NI2 and NI4 content over the territories subjected to natural and technogenic effects, reception of sufficient data on seasonal dependencies of tropospheric parameters being studied (in the Second Phase). Equipment used: Rusalka monoblock, Nikon D2X(s) digital photo camera; AF VR Nikkor ED 80-400f/4.5-5.6D lens with ultraviolet filter, bracket for attachment to the window, and Rusalka-Accessories set. Support hardware: Device TIUS /DKShG/PNSK, Laptop RSK1, and Software Package loading disk.]
Kotov & Suraev worked in the SM (Service Module), installing a new BRI Smart Switch Router computer, ASP Network Connection Adapter, new cables from the BRI to the old BRI location behind the medical locker and from the new BRI to the RSS1 and RSS2 laptops, plus from the ASP to future payload Ethernet connections behind panel 225.
Oleg downlinked the video recorded recently in the FGB.
In the SM, the FE-4 did the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS). [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]
Kotov also conducted the periodic checkout & performance verification of IP-1 airflow sensors in the various RS hatchways. [Inspected IP-1s are in the passageways PrK (SM Transfer Tunnel)–RO (SM Working Compartment), PkhO (SM Transfer Compartment)–RO, PkhO–DC1, PkhO–FGB PGO, PkhO-MRM2, FGB PGO–FGB GA, and FGB GA–Node-1].
FE-6 Creamer performed a water sample analysis using the TOCA (Total Organic Carbon Analyzer). Since the PWD (Potable Water Dispenser) and WRS was not available, TJ executed the run using a TOCA calibration bag,- not for a TOCA calibration, however, but for a nominal sample bag analysis.
FE-5 Noguchi and CDR “Zambo” Zamka worked in the US A/L (Airlock), resizing one of the spacesuits (EMU/Extravehicular Mobility Unit, #3018) for Robert Behnken, since EMU #3004, originally intended for Bob, had suffered a failure of its power harness, depriving the suit of the helmet camera and glove heater.
In the Kibo JPM (JEM Pressurized Module), Soichi installed SODI/IVIDIL (Selectable Optical Diagnostics Instrument / Influence of Vibration on Diffusion in Liquids) hardware, to be assembled to the SODI bottom plate in the MSG WV (Microgravity Science Glovebox Work Volume).
TJ performed the weekly 10-min. CWC inventory as part of on-going WRM (Water Recovery & Management) assessment of onboard water supplies. Updated “cue cards” based on the crew’s water calldowns are sent up every other week. [The current card (22-0003J) lists 77 CWCs (1,853.5 L total) for the five types of water now identified on board: 1. technical water (14 CWCs with 507.0 L, for Elektron electrolysis, incl. 157.3 L in 6 bags containing Wautersia bacteria, 134.2 L in 3 clean bags for contingency use, 215.5 L in 5 bags still requiring sample analysis, 2. potable water (9 CWCs with 366.7 L, of which 1 bag with 23.0 L contains Wautersia, 1 bag with 43.6 L requires sample analysis & 129.3 L in 3 bags are good for contingency use, 3. iodinated water (47 CWCs with ~881.1 L), 4. condensate water (1 bag with 28.1L [known leaker], 1 empty CWC, 3 bags with 50.4 L) and 5. waste/EMU dump and other (1 CWC with 20.2 L, 1 empty CWC). Wautersia bacteria are typical water-borne microorganisms that have been seen previously in ISS water sources. These isolates pose no threat to human health.]
Creamer activated/re-initialized ABRS (Advanced Biological Research System) in support of APEX-Cambium Payload Run 3 experiment operations. [TJ performed a flush of the payload cooling loop and performed replant 3A, retrieving appropriate TAGES (Transgenic Arabidopsis Gene Expression System) plant samples/plates from stowage for setting up the plants and starting their grow-out for Run 3.]
Noguchi configured the jumpers for providing O2
(oxygen) from the Shuttle directly to the US A/L for crew prebreathing during today’s first spacewalk campout in “Quest”.
At ~4:09am-5:09am EST this morning, the joint crew held an in-depth one-hour review of procedures for the EVA-1 spacewalk, with egress scheduled tonight at ~9:14pm EST.
At ~6:39am, Behnken (EV1) & Patrick (EV2) began their “campout” (nachalo desaturatsiy = desaturation start) in the A/L with hatch closure and depressurization of the CL (Crewlock) from 14.7 to 10.2 psi, followed by mask prebreathe (~6:39am-7:44am) and sleep from 8:14am-4:14pm tonight. A hygiene break is scheduled for 4:54pm-6:04pm. [Sleep for the ISS crew began at 7:44am.]
The crewmembers worked out with their regular 2-hr physical exercise on the TVIS treadmill (FE-1, FE-4), ARED advanced resistive exerciser (CDR, FE-4, FE-5, FE-6), T2/COLBERT advanced treadmill (CDR, FE-5, FE-6), and VELO bike ergometer with bungee cord load trainer (FE-1).
- Sleep shifting started with the Progress docking on 2/4. On 2/6, crew wake shifted earlier, to 5:40pm EST. 20A Undock will drive Crew Wake one and a half hours earlier to 4:09pm by FD12. This shift is accomplished by moving Crew Sleep 30 min earlier on FDs 6-10, and then again on FD11 and FD12. Wake/Sleep table:
|FD ||Cal ||GMT ||WAKE EST |
|SLEEP EST |
|4 ||11-Feb ||42 ||4:14pm (2/10) ||7:44am |
|5 ||12-Feb ||43 ||4:14pm (2/11) ||7:44am |
|6 ||13-Feb ||44 ||4:14pm (2/12) ||7:44am |
|7 ||14-Feb ||45 ||4:14pm (2/13) ||7:44am |
|8 ||15-Feb ||46 ||4:14pm (2/14) ||7:44am |
|9 ||16-Feb ||47 ||4:14pm (2/15) ||7:44am |
|10 ||17-Feb ||48 ||4:14pm (2/16) ||7:44am |
|11 ||18-Feb ||49 ||4:14pm (2/17) ||6:44am |
|12 ||19-Feb ||50 ||3:14pm (2/18) ||6:44am |
|13 ||20-Feb ||51 ||1:00am ||4:30pm |
No CEO (Crew Earth Observation) photo targets uplinked for today. Significant Events Ahead (all dates Eastern Time and subject to change)
02/10/10 -- STS-130/Endeavour/20A docking (12:57am)
- 02/12/10 -- EVA-1 (9:14pm; Node-3 install, OTP remove/stow)
- 02/12/10 -- Node-3 transfer
- 02/13/10 -- EVA-2 (9:14pm; Node-3 outfitting), Node-3 activation
- 02/15/10 -- Cupola relocation
- 02/16/10 -- EVA-3 (9:14pm; Node-3, Cupola, and PMA3 tasks)
02/17/10 -- Final cargo transfers, hatch closure, reboost
02/18/10 -- STS-130/Endeavour/20A undock (7:35pm)
02/20/10 -- STS-130/Endeavour/20A KSC deorbit burn (8:59pm)
02/20/10 -- STS-130/Endeavour/20A KSC landing (10:01pm)
03/18/10 -- Soyuz TMA-16/20S undock/landing --------------Three-crew operations-------------
03/18/10 -- STS-131/Discovery/19A – MPLM(P), LMC (launch ~1:30pm EST)
04/02/10 -- Soyuz TMA-18/22S launch – Skvortsov (CDR-24)/Caldwell/Kornienko
04/04/10 -- Soyuz TMA-18/22S docking --------------Six-crew operations-----------------
04/27/10 -- Progress M-03M/35P undock
04/28/10 -- Progress M-05M/37P launch
04/30/10 -- Progress M-05M/37P docking
05/14/10 -- STS-132/Atlantis/ULF4 – ICC-VLD, MRM-1 (~2:00pm EST)
05/10/10 -- Progress M-04M/36P undock
05/31/10 -- Soyuz TMA-17/21S undock/landing --------------Three-crew operations-------------
06/14/10 -- Soyuz TMA-19/23S launch – Wheelock (CDR-25)/Walker/Yurchikhin
06/16/10 -- Soyuz TMA-19/23S docking --------------Six-crew operations-----------------
07/xx/10 -- US EVA-15
07/xx/10 -- Russian EVA-25
06/28/10 -- Progress M-06M/38P launch
07/02/10 -- Progress M-06M/38P docking
07/26/10 -- Progress M-05M/37P undock
07/27/10 -- Progress M-07M/39P launch
07/29/10 -- Progress M-07M/39P docking
07/29/10 -- STS-134/Endeavour (ULF6 – ELC3, AMS-02) (~7:30am EST)
08/30/10 -- Progress M-06M/38P undock
08/31/10 -- Progress M-08M/40P launch
09/02/10 -- Progress M-08M/40P docking
09/15/10 -- Soyuz TMA-18/22S undock/landing
09/16/10 -- STS-133/Discovery (ULF5 – ELC4, PMM) (~12:01pm EST)
09/18/10 -- STS-133/Discovery (ULF5 – ELC4, PMM) docking
09/22/10 -- STS-133/Discovery (ULF5 – ELC4, PMM) undock
09/30/10 -- Soyuz TMA-20/24S launch – Kelly (CDR-26)/Kaleri/Skripochka
10/xx/10 -- Russian EVA-26
10/26/10 -- Progress M-07M/39P undock
10/27/10 -- Progress M-09M/41P launch
10/29/10 -- Progress M-09M/41P docking
11/15/10 -- Soyuz TMA-19/23S undock/landing
11/30/10 -- Soyuz TMA-21/25S launch – Kondratyev (CDR-27)/Coleman/Nespoli
12/15/10 -- Progress M-08M/40P undock
02/08/11 -- Progress M-09M/41P undock
02/09/11 -- Progress M-10M/42P launch
02/11/11 -- Progress M-10M/42P docking
03/30/11 -- Soyuz TMA-22/26S launch
xx/xx/11 -- Progress M-11M/43P launch
05/30/11 -- Soyuz TMA-23/27S launch
12/??/11 -- 3R Multipurpose Laboratory Module (MLM) w/ERA – on Proton.